OpenStudy (anonymous):

〖sin〗^(-1) (sin⁡(-3π/10) ). Is this straight forward and the answer -.122 and you use the calculator in radians?

OpenStudy (anonymous):

\[f^{-1}(f(x))=x\]

OpenStudy (anonymous):

well, not always with arcsine and sine, but in this case since \(-\frac{3\pi}{12}\) is in the interval \([-\frac{\pi}{2},\frac{\pi}{2}]\) you have no worries

OpenStudy (anonymous):

So it's just 3pie/10